OpenStudy (anonymous):

Use substition. 2m+n=1 m-n=8

OpenStudy (solomonzelman):

2m+n=1 m-n=8 ----> m=8+n 2(8+n)+n=1 16+2n+n=1 3n=-15 n=-15 plug that in into one of the equations m-(-15)=8 m+15=8 m=-7

OpenStudy (anonymous):

The answer should be (3,-5). Well at least that's what it says in the back of the book.

OpenStudy (anonymous):

@SolomonZelman 3n = -15 n=-5

OpenStudy (solomonzelman):

Yeah, sorry for this silly mistake!

OpenStudy (solomonzelman):

My 50th day, I am kind of new to the website...